20151201 » History » Revision 1

Revision 1/33 | Next »
Boyd, Steven, 30 November 2015 07:49

Commissioning and Ops Scheduling meeting : 1/12/2015

To be held in CR10 in the ATLAS Building @ 15:30 pm

Phone access : ReadyTalk number
US 866-740-1260
UK 0800-496-0576
Swiss 0800-70-53-51
Italy 800182592
Access Code: 2155402


  • Review of tasks and activities identified at last meeting
    • Absorber leak testing and cooldown plans
    • Diffuser leak fix
    • Decay Solenoid PSU Repair
    • SS Power SUpply and vacuum work
    • Control room systems test
  • Upcoming activities : chaired by Victoria
    • ISIS maintenance day tomorrow (Wed 25th)
      • Q4 cooling system flush
    • FC Soak test plan
    • Hydrogen/Vacuum system installation and commissioning
    • Hydrogen system detector system service (Dec 3rd)
  • Operations update
    • Progress on expert led run (Today)
    • Online software change control
  • AOB
    • To-do list for Maintenance shutdown (January 2016)


Present: Victoria, Colin, Craig, Mike, Melissa, Jaroslaw, Phil, Henry, Andy, Victoria, Josef, Mark, John G, Paul S, Pierrick, Chris R

  • Review of tasks
    • Absorber cooldown : found a leak in the absorber. Still under investigation so don't know how it impacts on cooldown plan yet.
    • Diffuser leak fix : went well. New diffuser went in but patch panel was leaking. New diffuser installed but still leaks - not as
      bad as previously. Can keep pressure there. John Govans wants to replace the He bottle - leak is still 4 bottles a day. Geoff has
      some ideas on how to get around it. Thinks main problem is pipe impedance - Melissa will meeting with Geoff and Colin to come up
      with a further plan.
      Andy: mustn't have situation when thin window is met by full bottle pressure since it could burst the window
    • Decay Solenoid PSU repair plan : PSU has been craned down, is being crated and will be sent back to FUG. FUG will look at it
      and decide what failed and why. Timescale: at least New Year.
      Parallel plan : consider putting together a power supply with controller in a relatively short timescale. DL looking into this.
      This was discussed at the MIPO and timescale is long. Need full logging and full quench protection guaranteed.
      Wait until FUG makes its assessment. Will test solenoid itself with a small power supply.
    • SS power supply and vacuum work : did half of it yesterday and the rest tomorrow. Next week Phill will be down to integrate
      it into the control system.
    • Control room systems test went OK.
  • Upcoming activities
  • Test run is fine
  • Wednesday : John can get in the DSA to do the cooling system flush on Q4 then need to check that the magnets are working.
    Temperature readouts from Q5/6 swapped. Need to check tomorrow or just swap in software.
    Pierrick - want to make certain sensors in the DSA are working. John Glover might be available to do this if he has time.
    Hall is open and Mark will do vacuum work
  • FC soak test plan (Thursday/Friday) : start powering FC, see if it gets to full current and then soak test : Steve Watson
    Need to put FC signal on the screen in the MCR. Work on that tomorrow. Pierrick can send a quench signal to MCR.
    Check with Paolo.
    Need to know how to cancel the search in an emergency - show them how to do this. Talk to Alan?
  • Hydrogen/Vacuum system installation and commissioning : discussed - leak checking
  • H detector system service - morning of Dec 3rd / can do a control room start-up using pulsers
  • Operations : Daresbury want access Dec 2,3,4 and Dec 8,9,10 - doesn't conflict with running
    Expert led run OK
    If anything has changed make sure the MOM knows. Talk with M. Courthold.

Updated by Boyd, Steven over 8 years ago · 1 revisions